OpenStudy (anonymous):

Find the vertex of the quadratic equation: y = 2x2 + 8x − 4.

OpenStudy (anonymous):

divide by two to get all that junk out

OpenStudy (anonymous):

x=-1 y =-4

OpenStudy (across):

You can use calculus: \[f(x)=2x^2+8x-4\]\[f'(x)=4x+8=0\]\[x=-2\]\[f(-2)=y=-12\]

OpenStudy (anonymous):

vertex = -b/2a where y=ax^2 + bx + c, so in this case: x=-8/2(2) = -2

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

another way would be to use\[(x-h)^2=4p(y-k)\]